Two 20.0-g ice cubes at –16.0 °C are placed into 265 g of water at 25.0 °C. Assuming no energy is transferred to or from the surroundings, calculate the final temperature of the water after all the ice melts.
Answer #1

specific heat of ice = 2.05 J/(g*C)

heat of fusion = 334 J/g

release heat when ice melts = (20 + 20)*2.05*16 + (20+20)*334 = 14672 J

now 40 g water is at 0 C and 265 g water is 25 C and 14672 J heat enegy is released.

14672 + 40*4.186*(T-0) = 265*4.186*(25 - T)

14672 = 27732.25 - 1109.29*T - 167.44*T

T = (27732.25 - 14672)/(1109.29 + 167.44)

T = 10.23 degree C [final temperatue]